10) If an investment earns 8% compounded continuously, how much should you deposit now to have $20,000 six years from now?

can you hand write the steps and the answer please. please i need the answer
Answer #1

Given that,

Interest rate r = 8% compounded continuously,

amount required in t = 6 years is FV = $20000

For continuous compounded, present value in account can be calculated using formula

PV = FV*e^(-r*t)

So, amount invested now PV = 20000*e^(-0.08*6) = $12375.67